OpenStudy (anonymous):

A circular plate has a diameter of 6 inches. What is the exact area of the plate? A. 36π square inches B. 6π square inches C. 3π square inches D. 9π square inches

OpenStudy (anonymous):

@TheSmartOne @inowalst @iambatman

OpenStudy (inowalst):

@saifoo.khan Can you help?

OpenStudy (anonymous):

i think it is c

OpenStudy (anonymous):

the fact that the radius is half of the diameter

OpenStudy (anonymous):

what do u guys and and where did u go? does somebody mind helping!!!

OpenStudy (anonymous):

i think c

OpenStudy (anonymous):

i actually think d now

OpenStudy (anonymous):

yep, i am 99.99% sure it i sd, do u agree

TheSmartOne (thesmartone):

Area=\(\sf\Large\left(\frac{d}{2}\right)^2\times 3.14\)

OpenStudy (anonymous):

it is d, i mean

OpenStudy (anonymous):

ok

TheSmartOne (thesmartone):

Yes, D is correct :)

OpenStudy (anonymous):

thanks a million :P

TheSmartOne (thesmartone):

Anytime :)